OpenStudy (anonymous):

Suppose F is defined as f(x)=2x-3. Solve for x in the inequality below. I f(x) -5 I less than or equal to 0.0001

OpenStudy (anonymous):

oh no \[|2x-3-5|\leq 0.0001\]

OpenStudy (anonymous):

or just \[|2x-8|\leq 0.0001\] which means \(2x\) is within \(0.0001\) units of 8

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

i got\[x \le 4.00005\]

OpenStudy (anonymous):

and i know you have to do 2 equations so the other would one would be \[\left| 2x-8 \right|\ge -0.0001\]

OpenStudy (anonymous):

and this is where im stuck